7.0 g of CO is place in a rigid 5.0 L vessel equipped with an injection port at 10oC. What is the pressure of CO vessel? How many grams of H2 must be injected into the vessel so that the resulting partial pressure CO will be ½ of the pressure when CO alone was present in the vessel?



1
Expert's answer
2021-10-27T06:45:10-0400

PV= nRT

5×0.5= n×8.314×283

2.5= 2352.86n

n= 0.0011mol

Molar mass of CO = 44.01

= 44.01×0.0011

= 0.047g